You put $4,750 into a bank paying 1.4% simple interest per year. Suppose you don't take anything out of the account. How much do you have after two years?
Thank you,
Shelley

Answer by nerdybill(7008)

You put $4,750 into a bank paying 1.4% simple interest per year. Suppose you don't take anything out of the account. How much do you have after two years?
.
M = P( 1 + i )^n
.
M is the final amount including the principal
P is the principal amount
i is the rate of interest per year
n is the number of years invested
.
In your case:
P is 4750
i is .014
n i2 2
.
Plugging in our values into:
M = P( 1 + i )^n
M = 4750(1 + .014 )^2
M = 4750(1.014)^2
M = 4750(1.028196)
M = $4883.93